larry want 50 pounds of tea that he can sell for $5.14 per pound. if he has tea that sells for $7.80 per pound and tea that sells for $4.30 per pound, how much of each should he use?
Answer by richwmiller(17219)

a+b=50,
4.30*a+7.80b=50*5.14
a=38., b=12.